# Evaluating and Enhancing LLMs for Multi-turn Text-to-SQL with Multiple Question Types

### v1

@@ 3.6 Result of MMSQL Evaluation
As shown in Table 3, GPT-4 Turbo demonstrated exceptional performance with a TDEX score of 67.0, while Gemini-1.5 Flash scored 65.8. GPT-4 Turbo also achieved the highest RQS of 5.8, indicating its strong overall capabilities. Meanwhile, Llama3-70B showed comparable TDEX performance to GPT-3.5, scoring 62.8 and 64.1 respectively, highlighting the robust potential of closed-source models in zero-shot settings. Notably, Llama3-8B achieved a TDEX of 64.0, slightly surpassing GPT-3.5.
Regarding question type recognition, GPT-4 achieved the highest average F1 score of 68.2. Open-source models also performed commendably, with Llama3-8B attaining an average F1 score of 64.2, demonstrating their competitive edge. While closed-source models exhibit superior performance across many metrics, open-source models are rapidly catching up.
@@
As shown in Table 3, both open-source and closed-source models face significant challenges with unanswerable and ambiguous questions, with the latter being particularly difficult. For example, despite demonstrating strong overall performance, GPT-4 Turbo, which achieves the highest question type F1 score of 68.2, exhibits much lower precision and recall for unanswerable questions (56.9 and 38.4) and ambiguous questions (25.9 and 70.0) compared to other types. This pattern is also consistent across other models, highlighting a widespread difficulty in accurately distinguishing such question types. In terms of natural language response quality, as shown in Table 4, Gemini-1.5 Flash performs well on unanswerable questions with a score of 9.78, but significantly drops to 4.31 on ambiguous questions. Codellama-7B is the lowest-performing model overall, with an average score of 5.54, particularly struggling with ambiguous questions at a score of only 3.57. This trend underscores a common challenge all models face in interpreting and responding to ambiguous queries, indicating a need for further development in handling ambiguity.
Table 4: Response quality scores of models in handling different types of questions.
@@
As illustrated in Figure 2, several representative models exhibit a notable decline in execution accuracy when addressing ambiguous queries, compared to their performance with clear, answerable queries. For example, GPT-4 Turbo’s execution accuracy decreases from 51.0% with answerable queries to 41.3% when handling ambiguous ones. Similarly, GPT-3.5 Turbo and Llama3-70B show significant declines, with accuracy falling from 47.3% to 34.5% and from 47.4% to 34.7%, respectively. However, execution accuracy improves significantly when models detect ambiguity and prompt users for clarification. For instance, when ambiguous queries are clarified, GPT-3.5 Turbo’s query match accuracy rises from 34.5% to 49.0%. This suggests that generating SQL for ambiguous queries without clarification results in diminished performance. Therefore, it is crucial for models to not only generate SQL queries but also effectively communicate any ambiguities to users. By clarifying uncertainties and helping users understand them, models can assist in obtaining accurate information. By incorporating the clarification process across multiple interactions, models can formulate more precise SQL queries, enhancing text-to-SQL systems’ usability and effectiveness.
Figure 2: EX performance analysis of different Q&A scenarios in model outputs of MMSQL test set.
4 Methodology
@@
As illustrated in Figure 3, we introduce an LLM-based multi-agent collaborative framework that utilizes intelligent agents with distinct functionalities to enhance text-to-SQL parsing. This framework comprises a Question Detector agent, which identifies the type of question and determines the appropriate response strategy. Accompanying it is the Question Decomposer agent, responsible for generating SQL queries by breaking down complex questions into simpler sub-questions. Additionally, two auxiliary agents: the Schema Selector and the SQL Refiner aid in refining the process. The Schema Selector narrows down the database schema to focus on relevant tables, while the SQL Refiner uses external tools for SQL execution and feedback, ensuring the queries are accurate and effective. A detailed introduction of these agents will be presented.
4.1 Schema Selector
The Schema Selector is designed to identify the essential subset of a database schema necessary for answering a given question. By selectively focusing on the most relevant tables and columns, the Selector curtails the interference from extraneous data, enhancing operational accuracy. This retrieving capability is particularly pertinent in complex text-to-SQL systems, such as Business Intelligence (BI) platforms, which often interact with large databases containing numerous tables and columns [2, 38, 39]. In our framework, the Selector is activated to choose tables and columns relevant to the interaction record solely when the schema size exceeds a specific threshold; otherwise, the full schema is employed. This adaptive mechanism ensures that our framework maintains efficiency and effectiveness across different database scales and complexities.
4.2 Question Detector
For answerable questions, the system proceeds to the Question Decomposer for further processing. In the case of ambiguous questions, where the query cannot be precisely mapped to the database schema, the system explains the ambiguity and attempts to rewrite it into answerable forms. For example, as shown in Figure 3, the question ”What is the flight number of Delta Airlines?” (Q2) could be interpreted in multiple ways: it might refer to flights departing from APG, as inferred from the previous context (Q1), or it could pertain to all flights in the database. Instead of rejecting such ambiguous queries, the system attempts to rewrite them into answerable questions and provides multiple potential answers in the final response.
@@ 4.2 Question Detector
4.3 Question Decomposer
The Question Decomposer plays a critical role in the multi-agent framework by breaking down complex questions into simpler, manageable Sub-Questions $\{Q_{1},Q_{2},\ldots,Q_{|Q|}\}$. For each Sub-Question $Q_{i}$, it generates a corresponding sub-SQL $\text{S}_{j}$ based on $\text{S}_{<j}$, and the schema subset to address specific parts of the original query. This process involves applying chain-of-thought reasoning, allowing the system to progressively solve each sub-question. As shown in Fig. 3, the Question Decomposer takes an ambiguous question, ”What are the Delta Airlines flight numbers for flights departing from APG?”, and decomposes it into sub-questions like ”Retrieve airlines for flights from SourceAirport APG” and ”Retrieve flight numbers for Delta Airlines from those flights.” Each sub-question is then translated into a sub-SQL, collectively forming the final SQL output.
4.4 SQL Refiner
The SQL Refiner, as an integral auxiliary operation of the Question Decomposer, serves a critical role in ensuring the precision and reliability of the SQL queries generated by our multi-agent framework. This component is indispensable for the inspection and correction of the responses produced, particularly when addressing the complex demands of text-to-SQL tasks. Frequently utilized, the Refiner has demonstrated its effectiveness in enhancing the output to meet stringent accuracy requirements [5, 6, 31]. Drawing a parallel to its counterparts in software development, such as metaGPT [30] and ChatDev [40], where intelligent agents are tasked with overall architectural design, code writing, and comprehensive testing, the SQL Refiner plays a similar role in the domain of database interactions. It adeptly adjusts SQL queries to accommodate various datasets, database schemas, and SQL generation styles, thereby ensuring the accuracy and effectiveness of the generated queries.
5 Experiments
5.1 Experimental Setup
@@
We conduct experiments on three prominent models: GPT-4 Turbo, Gemini-1.5 Flash, and Llama3-70b, as delineated in Section 3.6. These models were selected as our baselines to underscore the comparative effectiveness of our methodology. The objective of our experiments was to conduct a meticulous analysis elucidating how the incorporation of the multi-agent framework augments the performance of these models across a spectrum of evaluation metrics.
5.2 Overall Performance
Table 5: Performance metrics comparison on the MMSQL test set, demonstrating the improvements achieved by integrating the multi-agent framework with different models. Specific improvements are indicated in parentheses.
Model TDEX EX RQS F1 Score
GPT-4 Turbo 67.0 70.0 5.80 68.2
Multi-Agent+GPT-4 Turbo 70.0(+3.0) 74.4(+4.4) 5.98(+0.18) 69.5(+1.3)
Gemini-1.5 Flash 65.8 70.0 4.03 59.3
Multi-Agent+Gemini-1.5 Flash 69.4(+3.6) 73.7(+3.7) 7.05(+2.02) 70.7(+11.4)
Llama3-70B 62.8 66.4 3.86 59.8
Multi-Agent+Llama3-70B 70.7(+7.9) 74.8(+8.4) 6.64(+2.78) 70.8(+11.0)
[…]
77footnotetext: The best values are marked in bold.
Table 5 illustrates the performance of our method and baseline models on the MMSQL test set, highlighting the enhancements achieved by integrating our multi-agent framework with different models. The results demonstrate the benefits of our approach, with significant improvements across all metrics for each model. For example, the TDEX score increases from 62.8 to 70.7 and the F1 Score improves from 59.8 to 70.8 for Llama3-70B, showcasing the robustness of our framework in generating accurate SQL queries and natural language responses. These improvements underscore the effectiveness of our multi-agent framework in bolstering the capabilities of existing models to handle complex text-to-SQL tasks on the MMSQL benchmark.
5.3 Ablation Study
In Table 6, we presents the findings from an ablation study on the MMSQL test set, examining the impact of removing individual components from the multi-agent framework when integrated with Gemini-1.5 Flash. The complete model achieves its best performance with scores of 69.4 in TDEX, 73.7 in EX, 7.05 in Average RQS, and an F1 Score of 70.7. The study reveals that the removal of any component results in a decrease in performance across these metrics.
Table 6: Ablation study examining the impact of each component within the multi-agent framework on the model’s performance on the MMSQL test set.
@@ 5.3 Ablation Study
Multi-Agent + Gemini-1.5 Flash 69.4 73.7 7.05 70.7
@@
6 Discussion
In this study, we introduced the MMSQL test suite, a pioneering benchmark specifically designed to evaluate LLMs across complex multi-turn and multi-type text-to-SQL tasks. Our comprehensive evaluation of eight popular LLMs revealed that even the most advanced models struggle significantly with MMSQL’s diverse question types, particularly with ambiguous or unanswerable queries. This difficulty underscores the urgent need for models that possess enhanced capabilities for intent understanding and uncertainty management, as directly responding to ambiguous queries with SQL can severely degrade overall performance. Our findings are detailed in Table 3, highlighting the impressive potential of open-source models. Notably, the Llama3 series demonstrates performance that matches or even exceeds that of closed-source models on several metrics. Additionally, our findings indicate that training smaller models, such as CodeLlama-7B, with the MMSQL training set leads to substantial performance improvements. These results suggest pathways for developing more reliable and robust text-to-SQL systems. Furthermore, we designed a multi-agent framework tailored for handling multi-type text-to-SQL problems, which distinguishes itself from previous approaches that opted for abstention in the face of ambiguous queries [10, 7]. By refining the existing text-to-SQL multi-agent framework [5, 6] and incorporating a Question Detector Agent, our framework attempts to resolve ambiguities by requesting clarifications from users. This innovative approach not only enhances the model’s ability to deal with complex queries but also significantly improves user interaction. These advancements represent a significant step forward in the evolution of text-to-SQL technologies, demonstrating notable progress across a range of evaluation metrics.
@@ 6 Discussion
Our study has highlighted advancements in using LLMs for text-to-SQL applications, yet it also reveals several limitations that point to opportunities for further development. Firstly, we believe the performance of the open-source LLMs used in our tests could be enhanced by fine-tuning larger open-sourced LLMs. Secondly, our multi-agent framework is currently based on closed-source LLMs. Exploring open-source, domain-specific LLMs and further decomposing and annotating the MMSQL training set could synergistically enhance our framework’s effectiveness. This approach would provide tailored training data for each agent, allowing for targeted fine-tuning that leads to better results. This dataset also does not incorporate domain-specific knowledge bases, which limits its effectiveness in specialized fields such as medicine, law, or finance. Addressing these issues by integrating domain-specific knowledge into the training process could improve the model’s ability to accurately process and respond to specialized queries.
7 Conclusion
@@
Existing text-to-SQL evaluation methods often focus on increasing the complexity of SQL queries to enhance model performance. However, these approaches frequently encounter significant challenges in handling conversational dynamics, which further impede their development. In response to these challenges, this paper introduces the MMSQL test suite and proposes a novel multi-agent framework. The MMSQL test suite provides a structured approach to assess the capabilities of LLMs and offers an environment with diverse question types from a multi-turn perspective. Our analysis and experiments based on MMSQL illuminate the current limitations of LLMs. The multi-agent framework adeptly manages diverse question types. Ultimately, our work contributes to the democratization of data sharing, enhances the utility and robustness of text-to-SQL systems, and provides practical guidance for achieving high performance in this domain.
[…]
Funding National Natural Science Foundation, China (No. 62172123), the Key Research and Development Program of Heilongjiang (Grant No. 2022ZX01A36).
[…]
Data availability The data availability is outlined in Appendix B.
[…]
Code availability The code used in this article is publicly available on GitHub together with instructions on how to use the code.
Declarations
[…]
Supplementary information
The MMSQL dataset, experimental results, and the code for generating the training set are available in the Supplementary Files.
[…]
Conflict of interest The Authors listed in this article declare that they have no conflict of interest.
[…]
Ethical approval We prioritize ethical considerations and data privacy in our work. All content and labels used for our data augmentation and manual annotation are sourced exclusively from publicly available datasets.
  • Author contribution
@@ Appendix A Implementation Details
In all experiments, we employed the same chatting format for each LLM across all experiments. We applied greedy decoding strategies across all LLMs during inference and evaluation to ensure reproducibility. Training settings include lora_rank set to 64, lora_alpha set to 16, and learning rate set to 5e-5. The experiments were performed on a server with an Intel(R) Xeon(R) Gold 6133 CPU @ 2.50GHz and 4 NVIDIA A800 80GB PCIe GPU.
Appendix B Availability of data and material
We acknowledge the use of the CoSQL [^8] CoSQL datasets from [17] https://yale-lily.github.io/cosql and SParC [^9] SParC datasets from [12] https://yale-lily.github.io/sparc datasets, both licensed under the CC BY-SA 4.0 license [^10] https://creativecommons.org/licenses/by-sa/4.0/. This license allows for the sharing and adaptation of the datasets, provided that appropriate credit is given to the original creators and any modifications are distributed under the same license.
The MMSQL repository [^11] MMSQL dataset and associated scripts are accessible at https://github.com/mcxiaoxiao/MMSQL provides access to the augmented datasets generated by QDA-SQL, testing scripts, and experimental records. The implementation of the QDA-SQL method is available in a separate repository [^12] QDA-SQL scripts are accessible at https://github.com/mcxiaoxiao/QDA-SQL.

### v2

@@ III-E Result of MMSQL Evaluation
As shown in Table III, GPT-4 Turbo demonstrated exceptional performance with a TDEX score of 67.0, while Gemini-1.5 Flash scored 65.8. GPT-4 Turbo also achieved the highest RQS of 5.8, indicating its strong overall capabilities. Meanwhile, Llama3-70B showed comparable TDEX performance to GPT-3.5, scoring 62.8 and 64.1 respectively, highlighting the robust potential of closed-source models in zero-shot settings. Notably, Llama3-8B achieved a TDEX of 64.0, slightly surpassing GPT-3.5.
Regarding question type recognition, as shown in Table V, GPT-4 achieved the highest average F1 score of 68.2. Open-source models also performed commendably, with Llama3-8B attaining an average F1 score of 64.2, demonstrating their competitive edge. While closed-source models exhibit superior performance across many metrics, open-source models are rapidly catching up.
@@
As shown in Table V, both open-source and closed-source models face significant challenges with unanswerable and ambiguous questions, with the latter being particularly difficult. For example, despite demonstrating strong overall performance, GPT-4 Turbo, which achieves the highest question type F1 score of 68.2, exhibits much lower precision and recall for unanswerable questions (56.9 and 38.4) and ambiguous questions (25.9 and 70.0) compared to other types. This pattern is also consistent across other models, highlighting a widespread difficulty in accurately distinguishing such question types.
TABLE IV: Response quality scores of models in handling different types of questions: Answerable (Ans.), Unanswerable (Una.), and Ambiguous (Amb.), along with the average score.
@@
In terms of natural language response quality, as shown in Table IV, Gemini-1.5 Flash performs well on unanswerable questions with a score of 9.78, but significantly drops to 4.31 on ambiguous questions. Codellama-7B is the lowest-performing model overall, with an average score of 5.54, particularly struggling with ambiguous questions at a score of only 3.57. This trend underscores a common challenge all models face in interpreting and responding to ambiguous queries, indicating a need for further development in handling ambiguity.
TABLE V: Detailed performance metrics of the models on the MMSQL test set, with all values presented in percentages, presenting precision and recall for specific question types, along with Macro-F1 scores.
Model Ans. Una. Amb. Imp.  F1
Prec Rec Prec Rec Prec Rec Prec Rec
GPT-4 Turbo 90.3 89.8 25.9 70.0 56.9 38.4 100.0 98.0 68.2
GPT-3.5 Turbo 88.8 89.2 16.0 65.0 64.3 20.9 100.0 96.0 61.1
Gemini-1.5 Flash 85.9 95.8 26.3 50.0 58.3 8.1 100.0 95.4 59.3
Llama3-70B 84.9 95.2 27.5 55.0 80.0 9.3 100.0 92.7 59.8
Llama3-8B 88.3 93.2 21.4 60.0 83.3 23.3 100.0 96.7 64.2
SQLCoder-8B 84.6 99.6 77.8 35.0 0.0 0.0 100.0 98.0 59.7
Codellama-7B 93.9 16.5 4.3 85.0 96.6 66.3 56.8 100.0 46.9
Mistral-7B-v0.2 82.1 57.6 4.7 55.0 100.0 50.0 79.1 77.5 55.3
@@
As illustrated in Figure 3, several representative models exhibit a notable decline in execution accuracy when addressing ambiguous queries, compared to their performance with clear, answerable queries. For example, GPT-4 Turbo’s execution accuracy decreases from 51.0% with answerable queries to 41.3% when handling ambiguous ones. Similarly, GPT-3.5 Turbo and Llama3-70B show significant declines, with accuracy falling from 47.3% to 34.5% and from 47.4% to 34.7%, respectively. However, execution accuracy improves significantly when models detect ambiguity and prompt users for clarification. For instance, when ambiguous queries are clarified, GPT-3.5 Turbo’s query match accuracy rises from 34.5% to 49.0%. This suggests that generating SQL for ambiguous queries without clarification results in diminished performance. Therefore, it is crucial to generate SQL queries and effectively communicate ambiguities to users. By clarifying uncertainties and helping users understand them, models can assist in obtaining accurate information. Models can formulate more precise SQL queries by incorporating the clarification process across multiple interactions, enhancing text-to-SQL systems’ usability and effectiveness.
Figure 3: Performance analysis of model outputs on different Q&A scenarios in the MMSQL test set
IV Our Solution: LLM-based Multi-agent Collaborative Framework
IV-A Schema Selector
The Schema Selector is designed to identify the essential subset of a database schema necessary for answering a given question. By selectively focusing on the most relevant tables and columns, the Selector curtails the interference from extraneous data, enhancing operational accuracy. This retrieving capability is particularly pertinent in complex text-to-SQL systems, such as Business Intelligence (BI) platforms, which often interact with large databases containing numerous tables and columns. In our framework, the Selector is activated to choose tables and columns relevant to the interaction record solely when the schema size exceeds a specific threshold; otherwise, the full schema is employed. This adaptive mechanism ensures that our framework maintains efficiency and effectiveness across different database scales and complexities.
IV-B Question Detector
@@
For answerable questions, the system proceeds to the Question Decomposer for further processing. In the case of ambiguous questions, where the query cannot be precisely mapped to the database schema, the system explains the ambiguity and attempts to rewrite it into answerable forms. For example, as shown in Figure 4, the question "What is the flight number of Delta Airlines?" (Q2) could be interpreted in multiple ways: it might refer to flights departing from APG, as inferred from the previous context (Q1), or it could pertain to all flights in the database. Instead of rejecting such ambiguous queries, the system attempts to rewrite them into answerable questions and provides multiple potential answers in the final response.
@@
IV-C Question Decomposer
The Question Decomposer plays a critical role in the multi-agent framework by breaking down complex questions into simpler, manageable Sub-Questions $\{Q_{1},Q_{2},\ldots,Q_{|Q|}\}$. For each Sub-Question $Q_{i}$, it generates a corresponding sub-SQL $\text{S}_{j}$ based on $\text{S}_{<j}$, and the schema subset to address specific parts of the original query. This process involves applying chain-of-thought reasoning, allowing the system to progressively solve each sub-question. As shown in Fig. 4, the Question Decomposer takes an ambiguous question, "What are the Delta Airlines flight numbers for flights departing from APG?", and decomposes it into sub-questions like "Retrieve airlines for flights from SourceAirport APG" and "Retrieve flight numbers for Delta Airlines from those flights." Each sub-question is then translated into a sub-SQL, collectively forming the final SQL output.
IV-D SQL Refiner
The SQL Refiner, as an integral auxiliary operation of the Question Decomposer, serves a critical role in ensuring the precision and reliability of the SQL queries generated by our multi-agent framework. This component is indispensable for the inspection and correction of the responses produced, particularly when addressing the complex demands of text-to-SQL tasks. Frequently utilized, the Refiner has demonstrated its effectiveness in enhancing the output to meet stringent accuracy requirements [5]. Drawing a parallel to its counterparts in software development, such as metaGPT [32] and ChatDev [39], where intelligent agents are tasked with overall architectural design, code writing, and comprehensive testing, the SQL Refiner plays a similar role in the domain of database interactions. It adeptly adjusts SQL queries to accommodate various datasets, database schemas, and SQL generation styles, thereby ensuring the accuracy and effectiveness of the generated queries.
TABLE VI: Performance metrics comparison on the MMSQL test set, demonstrating the improvements achieved by integrating the multi-agent framework with different models. Specific improvements are indicated in parentheses.
Model TDEX EX RQS F1 Score
GPT-4 Turbo
w/o Multi-Agent 67.0 70.0 5.8 68.2
w/ Multi-Agent 70.0(+3.0) 74.4(+4.4) 6.0(+0.2) 69.5(+1.3)
Gemini-1.5 Flash
w/o Multi-Agent 65.8 70.0 4.0 59.3
w/ Multi-Agent 69.4(+3.6) 73.7(+3.7) 7.0(+3.0) 70.7(+11.4)
Llama3-70B
w/o Multi-Agent 62.8 66.4 3.9 59.8
w/ Multi-Agent 70.7(+7.9) 74.8(+8.4) 6.6(+2.7) 70.8(+11.0)
77footnotetext: The best values are marked in bold.
V Experiments
V-A Experimental Setup
@@
We conduct experiments on three prominent models in Section III-E: GPT-4 Turbo, Gemini-1.5 Flash, and Llama3-70b. These models were selected as our baselines to underscore the comparative effectiveness of our methodology. The objective of our experiments was to conduct a meticulous analysis elucidating how the incorporation of the multi-agent framework augments the performance of these models across a spectrum of evaluation metrics. For further details on the experimental implementation, refer to Appendix A.
V-B Overall Performance
Table VI illustrates the performance of our method and baseline models on the MMSQL test set, highlighting the enhancements achieved by integrating our multi-agent framework with different models. The results demonstrate the benefits of our approach, with significant improvements across all metrics for each model. For example, the TDEX score increases from 62.8 to 70.7 and the F1 Score improves from 59.8 to 70.8 for Llama3-70B, showcasing the robustness of our framework in generating accurate SQL queries and natural language responses. These improvements underscore the effectiveness of our multi-agent framework in bolstering the capabilities of existing models to handle complex text-to-SQL tasks on the MMSQL benchmark.
V-C Ablation Study
In Table VII, we presents the findings from an ablation study on the MMSQL test set, examining the impact of removing individual components from the multi-agent framework when integrated with Gemini-1.5 Flash. The complete model achieves its best performance with scores of 69.4 in TDEX, 73.7 in EX, 7.05 in Average RQS, and an F1 Score of 70.7. The study reveals that the removal of any component results in a decrease in performance across these metrics.
TABLE VII: Ablation study examining the impact of each component within the multi-agent framework on the Gemini-1.5 Flash’s performance on the MMSQL test set.
@@ V-C Ablation Study
Multi-Agent 69.4 73.7 7.05 70.7
@@
VI Discussion
In this study, we introduced the MMSQL test suite, a benchmark for evaluating LLMs on complex multi-turn and multi-type text-to-SQL tasks. Evaluating eight LLMs, we found significant challenges with MMSQL’s diverse question types, particularly ambiguous or unanswerable queries. This underscores the need for models with better intent understanding and uncertainty management. Notably, open-source models like the Llama3 series matched or surpassed closed-source models on several metrics, suggesting avenues for developing more reliable text-to-SQL systems. We also developed a multi-agent framework for handling multi-type text-to-SQL problems, unlike previous approaches that abstained when facing ambiguous queries [10, 7]. By refining the existing multi-agent framework [6] and adding a Question Detector Agent, our framework resolves ambiguities by requesting user clarifications, improving query handling and user interaction.
@@ VI Discussion
Despite these advancements, limitations remain. The performance of open-source LLMs could be improved with larger models. Our multi-agent framework currently relies on closed-source LLMs. Exploring open-source, domain-specific LLMs and further annotating the MMSQL training set could enhance our framework’s effectiveness [40], providing tailored training data for each agent and allowing targeted fine-tuning for better results [41]. Additionally, integrating domain-specific knowledge bases could improve the model’s ability to handle queries in specialized fields.
VII Conclusion
@@
Current text-to-SQL evaluation methods often emphasize SQL complexity, yet struggle with conversational dynamics, limiting their effectiveness. To address these challenges, we introduce the MMSQL test suite and a novel multi-agent framework. MMSQL offers a structured evaluation environment with diverse, multi-turn question types, highlighting limitations in existing LLMs. Our multi-agent framework effectively navigates these complexities, enhancing the robustness and usability of text-to-SQL systems. This work advances data accessibility and provides practical insights for achieving high performance in this domain.
@@ Appendix A Implementation Details
To ensure reproducibility, we applied greedy decoding strategies during both inference and evaluation. The experiments were conducted on a server equipped with an Intel Xeon Gold 6133 CPU and four NVIDIA A800 80GB PCIe GPUs.
TABLE VIII: Evaluation Criteria for Response Quality
Criterion Score Description
Relevance 0 The response is completely irrelevant.
1 The response is partially relevant but misses key details.
2 The response is fully relevant and adequately addresses the question.
Clarity 0 The response is incomprehensible.
1 The response is mostly clear with minor ambiguities.
2 The response is very clear and easy to understand.
Completeness 0 The response does not address the question at all.
1 The response covers most aspects of the question but lacks some details.
2 The response thoroughly addresses all aspects of the question.
Accuracy 0 The response contains factually incorrect information.
1 The response is partially accurate with some errors.
2 The response is completely accurate.
Utility 0 The response does not meet the user’s needs or explain the context of the question.
1 The response somewhat meets the user’s needs and provides partial explanations.
2 The response excellently meets the user’s needs and clearly explains the context or ambiguity of the question.
Appendix B Design and Effectiveness Study of QDA-SQL for Data Generation
In this study, we use QDA-SQL to expand the original datasets by inputting samples from SParC and CoSQL, while adhering to MMSQL’s classification and formatting standards. As depicted in Figure 5, QDA-SQL first generates multi-turn question-and-answer pairs in the Generate Interactions phase. During this phase, each Q&A pair is assigned a randomly selected thematic relation (as defined by [12]) and question type, referencing the existing dialogue history to maintain contextual relevance. In the subsequent Verify and Refine phase, Gemini Pro optimizes, scores, and filters these pairs to ensure alignment with our question type requirements and maintain high quality.
Figure 5: Overview of QDA-SQL processes.
To evaluate the quality of the data generated by QDA-SQL, A manual review of the automatically filtered results was conducted to evaluate the quality of the data generated by QDA-SQL, focusing on both classification accuracy and the quality of the annotated natural language question answering. A comparison of samples before and after the Verify and Refine process revealed that Gemini Pro successfully identified 94% (47 out of 50) of the misclassified samples. This finding underscores the effectiveness of the automatic filtering process in enhancing data quality and supporting high classification accuracy.
Additionally, an automatic evaluation framework based on GPT-4 was implemented, the effectiveness of which has been confirmed by numerous studies [42, 43]. Critical dimensions include completeness, relevance, and utility in answering the annotated natural language question. To address potential order bias [44], the placement of the original dataset and the enhanced dataset was alternated in pairwise comparisons, positioning the enhanced dataset first for odd IDs and second for even IDs. As illustrated in Figure 6, the enhanced dataset demonstrated superior performance, with an overall assessment indicating that 62% of the QDA-SQL enhanced dataset is considered superior to the manually annotated original SParC and CoSQL samples.
Figure 6: Pairwise comparison of original and QDA-SQL enhanced dataset annotation quality across different criteria.
Appendix C LLM-Assisted RQS Metric Evaluation and Cost
We used GPT-4o-mini to evaluate natural language responses across five dimensions: Utility, Accuracy, Completeness, Clarity, and Relevance (see Table VIII). Only the RQS metric in our MMSQL testing employed the OpenAI GPT-4o-mini API through LLM-assisted rating. We scored each of four metrics from 0 to 2, with a maximum total score of 10. This process incurred API costs—currently \$0.15 per million input tokens and \$0.60 per million output tokens for GPT-4o-mini. Scoring our 149 test rounds cost approximately \$0.036, significantly less than manual scoring, which relies heavily on human labor.